w varies directly as x and y inversely as v square and w=1200 when x=4, y=9 and v=6 find w when x=3 , y=12, and v=9
Answer by abdulaleem(11)

W = kXY/V²
Step 1: Find k; W = 1200 ; X=4; Y=9
1200 = k (4)(9)/(6)²
1200 = k(36)/36
k = 1200
Step 2: k = 1200; W = ? X=3; Y=12; V=9
W = kXY/V²
W = (1200)(3)(12)/(9)²
W = 43,200/81
W = 533 and 27/81 or 533 1/3